Given the Cobb - Douglas production q1 = 50l0.3K0.5,q2 = 50L0.4K0.6,and q3 =50l0.5K0.6..Calculate the level of output when L = 10 and K = 15 for each production function.. Calculate the level of output when both inputs double L = 20 and K =30...What types of returns to scale does each production function exhibit


1
Expert's answer
2022-03-09T10:55:45-0500

"Q_1=50L^{0.3}K^{0.5}"


"Q_2=50L^{0.4}K^{0.6}"


"Q_3=50L^{0.5}K^{0.6}"


Level of output When;

L=10 and K=15


"Q_1=50\times 10^{0.3}\times 15^{0.5}"

"=386.38"


"Q_1=50\times 10^{0.4}\times 15^{0.6}"


"=637.71"


"Q_1=50\times 10^{0.5}\times 15^{0.6}"


"=802.83"


The level of output when both inputs double L = 20 and K =30


"Q_1=50\times 20^{0.3}\times 30^{0.5}"


"=672.73"


"Q_2=50\times 20^{0.4}\times 30^{0.6}"


"=1,275.42"


"Q_3=50\times 20^{0.5}\times 30^{0.6}"


"=1,720.91"


 Returns to scale do each production function exhibit;


Q1- 0.8 is <1 Hence the production function exhibits decreasing return to scale.


Q2 - 1=1 Hence the production function exhibits decreasing return to scale.


Q3- 1.1 is >1 Hence the production function exhibits increasing return to scale.